Compute the mean and variance of random variable whose distribution is uniform on the interval[a,b].


(a) It is not enough to simplify state values. You must give the details of the computation

1
Expert's answer
2021-11-19T04:44:23-0500

Probability density function:


"f(x)=\begin{cases}
 \frac{1}{b-a} &\text{for } a\le x\le b \\
 0 &\text{for } x<a\ or\ x>b
\end{cases}"


mean:


"E(X)=\int^b_axf(x)dx=\int^b_a\frac{x}{b-a}dx=\frac{x^2}{2(b-a)}|^b_a=\frac{b^2-a^2}{2(b-a)}=\frac{b+a}{2}"


variance:

"V(X)=E(X^2)-(E(X))^2"


"E(X^2)=\int^b_ax^2f(x)dx=\int^b_a\frac{x^2}{b-a}dx=\frac{x^3}{3(b-a)}|^b_a=\frac{b^3-a^3}{3(b-a)}=\frac{a^2+ab+b^2}{3}"


"(E(X))^2=\frac{(b+a)^2}{4}"


"V(X)=\frac{a^2+ab+b^2}{3}-\frac{(b+a)^2}{4}=\frac{4(a^2+ab+b^2)-3(a^2-2ab-b^2)}{12}=\frac{a^2-2ab+b^2}{12}=\frac{(a-b)^2}{12}"